# Large Language Models Can Learn Temporal Reasoning

### v3

@@ 5 Related Work
Language-based Temporal Reasoning. Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023; Vashishtha et al., 2020). The vision here is to help LMs understand temporal concepts. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge. Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
Reasoning towards LMs. Although LLMs exhibit some emergent behaviors (Zhang et al., 2023a, 2024; Lai et al., 2024a, b; Wang et al., 2024; Chen et al., 2024a), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
Reasoning over Knowledge Graphs. Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Yang et al., 2017; Sun et al., 2018; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021; Huang et al., 2023; Chen et al., 2023, 2024b; Wan et al., 2024; Li et al., 2024). Moreover,  (Liu et al., 2022; Yang et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between language-based and symbolic-based temporal reasoning. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023b; Yuan et al., 2023), which limit their application to other tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to deliberate reasoning.
@@ 6 Conclusion
TG-LLM, a novel framework for LMs, has been proposed to improve their performance on temporal reasoning. To produce reliable final answers, our framework equips LLMs with the temporal graph and intermediate reasoning steps. Extensive experiments indicate that TG-LLM achieves better performance than existing pipelines. An interesting direction for future work is to extend it to more complex applications such as inductive and abductive reasoning. Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks.
@@
Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the quality of generated intermediate steps.
@@ Appendix A Dataset Statistics of TGQA
Figure 6: Number of facts distribution in TGQA.
Figure 7: Types of questions distribution in TGQA.
Appendix B Experiment Details
@@
Graph Data Augmentation.  We show with an example our graph data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on graph-based QAs. To contrust, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
@@ Appendix D Fine-grained Results of TGQA
We provide the fine-grained results of TGQA in Table 11. In TGQA, there are nine types of questions as explained in Appendix A. We show all the models with different strategies. Zero-shot performance is also considered to investigate the effect of in-context examples. Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results. For zero-shot learning results, the format of generation is not guaranteed. The original accuracy is very low since rule-based parser cannot handle it. Instead, we use GPT-3.5 as parser to extract the answer from the generation. Also, we only adopt one example in context due to context length limitation.
It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4. Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories. However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well. On the other hand, we show that our strategies improve the performance step by step. Note that our model achieves the best overall performance by improving all the categories. Specifically, deliberate reasoning with CoT bootstrapping brings better performance for most categories. Furthermore, graph data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.

### v4

@@ 5 Related Work
Language-based Temporal Reasoning. Recently, language-based TR has gained substantial research focus  (Liu et al., 2023, 2024b; Chen et al., 2024a, b; Wang et al., 2024; Cheng et al., 2024a; Xia et al., 2024). The vision here is to help LMs understand temporal concepts and logic such that they can perform more complicate tasks. Existing methods mainly solve this problem with time-aware language modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose specific pre-training/fine-tuning strategies for robust TR. On the other hand,  (Ning et al., 2019; Zhou et al., 2020, 2021; Yang et al., 2020, 2023a) design auxiliary objectives to introduce external temporal knowledge. Although these methods made some progress, representation learning for the underlying structure and logic of TR are either ignored or not explicitly involved.
@@
Reasoning towards LMs. Although LLM has exhibited some emergent behaviors (Zhang et al., 2024a; Lai et al., 2024a, b; Lyu et al., 2024), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is. Existing methods that try to elicit or enhance reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use in-context learning where some demonstrations including intermediate reasoning steps are provided in prompt.  (Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with the intermediate thinking process before producing final answers. On the other hand,  (Schick et al., 2023; Kynoch et al., 2023; Xu et al., 2023) combine LMs with domain-specific external tools, empowering the model to perform more complex tasks that require reasoning and interactions with environment.
Reasoning over Knowledge Graphs. Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Sun et al., 2018; Yang and Song, 2020; Huang et al., 2023; Wan et al., 2024; Li et al., 2024). Related work includes symbolic reasoning over temporal KGs  (Yang et al., 2022; Xiong et al., 2023, 2024), and language-based reasoning over static KGs  (Cheng et al., 2024b; Zhang et al., 2024b; Liu et al., 2024a; Zhao et al., 2024; Xu et al., 2024; Wei et al., 2024). In this paper, we build the connection between language-based and symbolic-based TR. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to graph reasoning.
@@ 6 Conclusion
TG-LLM, a novel framework for LMs, has been proposed to improve their performance on temporal reasoning. To produce reliable final answers, our framework equips LLMs with the temporal graph and intermediate reasoning steps. Extensive experiments indicate that TG-LLM achieves better performance than existing pipelines. An interesting direction for future work is to extend it to more complex applications such as inductive and abductive reasoning. Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks as well.
@@
Graph-augmented approaches  (Jin et al., 2024; Fan et al., 2024; Shang and Huang, 2024; He and Hooi, 2024) including TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Hao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the performance.
@@
Acknowledgements
This work was supported by a sponsored research award by Cisco Research.
@@ Appendix B Experiment Details
Graph Data Augmentation.  We show with an example our graph data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on graph-based QAs. To contrast, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
@@ Appendix D Fine-grained Results of TGQA
We provide the fine-grained results of TGQA in Table 11. In TGQA, there are nine types of questions as explained in Appendix A. We show all the models with different strategies. Zero-shot performance is also considered to investigate the effect of in-context examples. Note that we do not include zero-shot performance of the Llama2 family since they are not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results. For zero-shot learning results, the format of generation is not guaranteed. The original accuracy is very low since rule-based parser cannot handle it. Instead, we use GPT-3.5 as parser to extract the answer from the generation (might introduce errors). Also, we only adopt one example in context due to context length limitation (long-context LLM performance could increase with more in-context demonstrations).
It can be seen that in-context demonstrations are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4. Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories. However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well. On the other hand, we show that our strategies improve the performance in a progressive manner. Note that our model achieves the best overall performance by improving on all the categories. Specifically, deliberate reasoning with CoT bootstrapping brings better performance for most categories. Furthermore, graph data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.